The mean incubation time for a type of fertilized egg kept at 100.7 °F is 21days. Suppose that the incubation times are approximately normally distributed with a standard deviation of 11day.

(a) What is the probability that a randomly selected fertilized egg hatches in less than 20days?

(b) What is the probability that a randomly selected fertilized egg takes over 22days to hatch?

(c) What is the probability that a randomly selected fertilized egg hatches between 19 and 21days?

(d) Would it be unusual for an egg to hatch in less than 18days? Why?
Answer #1

Solution :

Given that ,

mean = = 21

standard deviation = = 11

a) P(x < 20 ) = P[(x - ) / < (20- 21) / 11 ]

= P(z < -0.09)

= 0.4641

probability = 0.4641

b)

P(x > 22 ) = 1 - p( x< 22)

=1- p P[(x - ) / < (22-21) /11 ]

=1- P(z < 0.09)

= 1 - 0.5359 = 0.4641

Probability =0.4641

c)

P( 19< x < 21) = P[(19-21)/11 ) < (x - ) /  < (21-21) /11 ) ]

= P(-0.18 < z < 0)

= P(z <0 ) - P(z < -0.18)

= 0.5 - 0.4286 = 0.0714

Probability = 0.0714

d)

P(x < 18 ) = P[(x - ) / < (18- 21) / 11 ]

= P(z < -0.27)

= 0.3940

probability = 0.3940

Would it be usual for an egg to hatch in less than 18 days ,

because probability greater than 0.05 .